1 1/3+2+3/4=
1/4divided by3/2=
ratio of 2 to 5 = what percent
12:30 as a percentage

Answer by stanbon(48535)

1 1/3+2+3/4 = (4/3) + 2 + (3/4) = [14 + 24 + 9]/12 = 35/12
---------------------
1/4divided by3/2 = (1/4)/(3/2) = (1/4)*(2/3) = 2/12 = 1/6
----------------------

ratio of 2 to 5 = what percent
2/5 = x/100
x = 100(2/5) = 40
2/5 = 40%
-----------------
12:30 as a percentage
12/30 = x/100
x = 100(12/30)
x = 100(4/10) = 40
12:30 = 40%
=============
Cheers,
Stan H.